OpenStudy (anonymous):

write 16^(-3/4) in radical notation

OpenStudy (anonymous):

the numerator is the power, the denominator is the root, and the minus sign means take the reciprocal you get \[\frac{1}{\sqrt[4]{16^3}}\] as a first step

OpenStudy (anonymous):

then the fourth root of 16 is 2, since \(2^2=16\) and \(2^3=8\) your "final answer" is \[\frac{1}{8}\]